# First, Learn What You Don’t Know: Active Information Gathering for Driving at the Limits of Handling

### v1

@@ Abstract
Combining data-driven models that adapt online and model predictive control (MPC) has enabled effective control of nonlinear systems. However, when deployed on unstable systems, online adaptation may not be fast enough to ensure reliable simultaneous learning and control. For example, controllers on a vehicle executing highly dynamic maneuvers may push the tires to their friction limits, destabilizing the vehicle and allowing modeling errors to quickly compound and cause a loss of control. In this work, we present a Bayesian meta-learning MPC framework. We propose an expressive vehicle dynamics model that leverages Bayesian last-layer meta-learning to enable rapid online adaptation. The model’s uncertainty estimates are used to guide informative data collection and quickly improve the model prior to deployment. Experiments on a Toyota Supra show that (i) the framework enables reliable control in dynamic drifting maneuvers, (ii) online adaptation alone may not suffice for zero-shot control of a vehicle at the edge of stability, and (iii) active data collection helps achieve reliable performance.
@@ I Introduction
Controlling unstable nonlinear systems remains challenging [1]. For example, guiding a vehicle through dynamic and unstable drifting maneuvers is a difficult task where slight decision-making mistakes can lead to spin outs or crashes. In recent years, advanced control methods have been developed with the goal of giving autonomous vehicles the ability to leverage their full handling potential [2, 3, 4, 5, 6], with possible applications to the future design of advanced driver-assistance systems and autonomous driving safety systems.
In this setting, data-driven model predictive control (MPC) methods have gained wide popularity [7] due to their high performance, capability to use expressive learned models, and ability to account for constraints such as actuator limits and obstacle avoidance. In particular, data-driven models that adapt using online data are able to reduce model errors and sim-to-real discrepancies at deployment time [8, 9, 10, 11]. However, simultaneously controlling and learning unstable systems is notoriously difficult. For instance, successfully executing a drifting maneuver requires a sufficiently-accurate initial model to precisely initiate the drift, and online adaptation alone may not enable identifying a model fast-enough to execute the maneuver, see Section V-C.
Autonomous driving at the limits, therefore, presents a challenge for data-driven MPC with online adaptation, due to its fast-paced nature and the inherent instability of the executed maneuvers. These difficulties raise the following questions: How can we effectively identify uncertain nonlinear dynamics to ensure that subsequent control is reliable? How should informative data be collected efficiently and safely to minimize time-intensive and potentially expensive testing on hardware?
Figure 1: We propose an active information gathering framework that guides data collection to quickly learn a model in a controlled (i.e., safe) setting. Adapting on this information-rich data enables reliable control in dynamic drifting trajectories. In contrast, directly using the model without adapting prior to deployment may lead to unreliable control performance such as spinning out.
@@
  • We present a data-driven vehicle dynamics model that combines a physics-based model and last-layer Bayesian meta-learning model. By learning expressive neural network features and only adapting the last layer, the model is conducive to rapid online adaptation and offers a structured and interpretable uncertainty representation.
  • We present an active information gathering method to identify an accurate model as fast as possible while satisfying constraints during data collection.
  • We extensively validate the proposed approach on a Toyota Supra in dynamic drifting maneuvers. In particular, results show that online adaptation of the meta-trained model may not be sufficient to reliably initiate a drift, and that active information gathering significantly helps quickly identifying a model and enabling robust drifting maneuvers.
From a vehicle control perspective, our results demonstrate that executing drifting maneuvers is possible by adapting a model on autonomously-generated non-drifting trajectories. Specifically, information-rich trajectories do not necessarily involve drifting, and adapting on such data aids in robust execution of challenging drifting maneuvers.
@@ II Related Work
Data-driven models are increasingly popular for control due to their ability to leverage expressive models and adapt them online given measurements of the system, see [7] and [12] for recent reviews. For example, model predictive control (MPC) approaches leveraging neural network models have been widely used in the recent years for dynamic driving at the limits of handling [9, 10, 11, 4]. These approaches adapt models online with the implicit assumption that the model will adapt fast enough during the task and enable satisfactory performance, which may be difficult for systems where small initial errors quickly amplify.
Online learning and control algorithms that actively seek informative data to improve the model [13, 14, 15, 16, 17, 18, 19, 20] often result in improved control performance compared to passively adapting a model while performing a task, and allow enforcing safety constraints during the data collection process [15, 20]. Leveraging such active learning approaches for controlling unstable nonlinear systems remains challenging. Indeed, the tradeoffs between seeking informative data (exploration) and performing the task (exploitation) are well-known; unstable systems are particularly difficult to simultaneously learn and control [21, 22]. These tradeoffs motivate the design of a framework where active learning is done prior to deployment in a controlled setting, with the objective of learning a model for control with minimal experimentation.
Our proposed method leverages a Bayesian modeling approach [23] to last-layer meta-learning [24, 25, 23, 26, 27], and combines it with model predictive control. We take inspiration from [20], and present a physics-informed data-driven vehicle model, learning the residual between a known nominal model and the true vehicle model of interest.
@@ III Bayesian vehicle dynamics model
Central to our approach is a vehicle dynamics model that is expressive, captures uncertainty, and enables rapid online adaptation. To this end, we combine a physics-based nominal bike model with a Bayesian last-layer meta-learning model. The former encodes prior knowledge about the vehicle, while the latter captures model mismatch and enables rapid adaptation over expressive neural network features. The vehicle state is expressed in curvilinear coordinates [5, 11], see Figure 2. As in [5, 11], we do not use brake inputs in this work. The states and control inputs of the vehicle are
@@
where $(n,m)=(6,2)$ are the state and control dimensions, $r$ is the yaw rate, $v$ is the total velocity, $\beta$ is the sideslip, $\omega_{r}$ is the rear wheelspeed, $e$ and $\Delta\varphi$ are the lateral and angle deviations to the reference trajectory, $\delta$ is the steering angle, and $\tau$ is the engine torque, respectively.
@@
where $k\in\mathbb{N}$ and $\epsilon_{k}\sim\mathcal{N}(0,\text{diag}(\sigma_{1}^{2},\dots,\sigma_{n}^{2}))$ are Gaussian-distributed disturbances that are independent over time $k$ and state dimensions, $i$. $\xi$ represent unknown parameters of the dynamics (e.g., corresponding to unmodeled phenomena) and $\tilde{u}_{k}=(u_{k},\dot{u}_{k})$ includes the control input rate $\dot{u}_{k}=(\dot{\delta}_{k},\dot{\tau}_{k})$ at time $k$. Including $\dot{u}_{k}$ in $\tilde{u}_{k}$ and (2) accounts for the fact that MPC plans $(u_{0},u_{1},\dots,u_{N})$ are executed by a low-level controller via linear interpolation on the control horizon. That is, $u(k\Delta t+\tau)=u_{k}+\tau\dot{u}_{k}$ with $\Delta t=0.1\mathrm{s}$ and $\dot{u}_{k}=(u_{k+1}-u_{k})/\Delta t$ for $\tau\in[0,\Delta t)$.
@@
$x_{k+1}=h(x_{k},u_{k})+g_{\theta}(z_{k})+\epsilon_{k},$ (3)
where $h:^{n}\times^{m}\to^{n}$ is given by Euler integration of a standard bike model [11], and the residual $g_{\theta}$ is modeled by $g_{\theta}(z)=(g_{1},\dots,g_{n})(z)$ with inputs
@@
and each dimension $i$ of $g_{\theta}(\cdot)$ is linear over learned features $\phi_{i}(z)\in^{d}$ for some parameters $\theta_{i}\in^{d}$, that is
@@
We omit the variables $e$ and $\Delta\varphi$ since they do not affect the dynamics of any of the other state variables and their dynamics can be explicitly computed from geometry, see [5]. That is, given the state in (1), we let $g_{5}=g_{6}=0$.
The features $\phi_{i}(\cdot)$ are neural networks that are learned jointly with the parameters $\theta$ via last-layer Bayesian meta-learning [23, 28, 20] to yield an expressive model capable of rapid adaptation with uncertainty estimates. Specifically, viewing $g_{i}(z_{k})$ as a neural network, the last-layer parameter $\theta_{i}$ are parameterized using a Gaussian distribution
@@
with mean parameters $\bar{\theta}_{i}$ and positive definite precision matrices $\Lambda_{i}$. Following [20], we model independent Gaussian distributions of each $\theta_{i}$, such that the joint distribution for the last layer parameters $(\theta_{1},\dots,\theta_{n})$ is Gaussian.
Computing the posterior: The linear structure of $g_{i}(z)$ enables the computation of the last-layer parameters $\theta_{i}$ via Bayesian linear regression. Specifically, given prior distributions $\mathcal{N}(\overline{\theta}_{i,0},\sigma_{i}^{2}\Lambda_{i,0}^{-1})$ for $\theta_{i}$, and a state-control trajectory $\mathcal{D}_{T}=\{(z_{k},x_{k+1})\}_{k=0}^{T}$, the posterior $\theta_{i,T}|\mathcal{D}_{T}$ is Gaussian and can be computed recursively from $Q_{i,0}=\Lambda_{i,0}\bar{\theta}_{i,0}$ by
@@
With the proposed model, the mismatch between the true dynamics $f$ and the nominal model $h$ through $g_{\theta}$ is learned such that the one-step ahead predictions are Gaussian by design, facilitating both meta-learning and information gathering.
Offline Bayesian meta-training: The prior parameters $(\bar{\theta}_{i,0},\Lambda_{i,0}^{-1})$, features $\phi_{i}$, and noise covariances, $\sigma_{i}^{2}$, are pre-trained offline to obtain an expressive model that can be adapted online using the update rule (6). Specifically, given a dataset of $J$ trajectories corresponding to different parameters $\xi$ of the true system (2) and a meta-training horizon $T$, we minimize the negative posterior log-likelihood objective
@@
As a dataset, we train on both successful and unsuccessful manual and autonomous drifting data from vehicle sensor data, totalling approximately 19 minutes of driving. This data corresponds to $35$ different trajectories, i.e., realizations of $f_{\xi}$, and contains a total of $11382$ data points. Following [11], we additionally augment our dataset with a mirrored set of data by negating the signs of the yaw rate, sideslip, and steering angle to simulate turning in the opposite direction. This augmentation doubles our dataset size.
For our architecture, we use a $\tanh$-feedforward neural network for each $\phi_{i}$ with two hidden layers of width $128$ and the dimension of each $\theta_{i}$ is $16$. Following [20], we share weights and biases across each $\phi_{i}$ except for their last layer weights and biases $(W_{\text{out}}^{i},b_{\text{out}}^{i})\in^{16\times 128}\times^{16}$. We use a meta-training horizon $T=10$ and train our model using Adam with an initial learning rate of $10^{-3}$ and an exponential decay rate of $0.9975$ over $1000$ epochs.
@@ IV Information Gathering MPC Framework
We present an active information gathering framework, shown in Figure 3, to identify nonlinear dynamics as quickly as possible for reliable subsequent deployment within MPC. First, offline, we meta-train a Bayesian vehicle model (Section III) using a dataset of vehicle trajectories. Second, we leverage the model’s expressive feature space $\phi$ and uncertainty representation encoded by $(\bar{\theta},\Lambda^{-1})$ to safely collect informative data and refine the model. After information-guided adaptation, we leverage the refined vehicle model within an MPC framework to control the vehicle in challenging driving scenarios such as drifting. Next, we describe the formulations used for MPC and information gathering.
@@
We define an MPC formulation that enables tracking dynamic trajectories such as drifting maneuvers parameterized by a reference trajectory $x_{\mathrm{ref}}$. To favor smooth control inputs, we penalize fast changes in the control inputs, and define
@@
$f_{\bar{\theta}}(x_{k},u_{k},u_{k+1})=h(x_{k},u_{k})+g_{\bar{\theta}}(z_{k}),$
where $g_{i}(z_{k})=\bar{\theta}_{i}^{\top}\phi_{i}(z_{k})$ for each $i=1,\dots,n$.
Directly deploying an MPC controller that recursively solves OCP using the prior model parameters $\bar{\theta}_{i,0}$ after offline meta-training may lead to poor performance in practice, even if one adapts the model online during the task, see Section V. These observations motivate an active data collection approach to refine the model prior to deployment in challenging applications. Since data collection on a vehicle may be time-consuming and expensive, this raises the question: What data should we collect to best identify the unknown dynamics?
The key insight in answering this question is that learning the model as quickly as possible amounts to maximizing the information gain from future observations, i.e., to maximizing the mutual information between observations and the true system (see [29, Chapter 2] and [30, Chapter 8]). Thus, to steer the system towards regions of high model uncertainty, we define the information gathering objective
@@
where $\alpha>0$ weighs the information gain compared to the nominal MPC objective. The nominal cost $\ell$ encourages smooth control inputs and the constraints (8b)-(8e) ensure that the vehicle remains on the track during data collection.
The information gathering objective (9) is derived in [20] by leveraging the last-layer uncertainty representation of the model $g_{\theta}$, which yields a closed-form expression of the one-step mutual information between an observation and the dynamics. Similar objectives are used for active learning in [31, 32, 16]. As discussed in [20], $\ell_{\mathrm{info}}$ is an approximation of the true total information gain over a trajectory, which is not exactly the sum of the expected information gains per timestep. We found that incorporating the linear regression updates (6) in the objective (by defining the cost ~$\tilde{\ell}_{\text{info}}=\frac{1}{2}\sum_{k=0}^{N-1}\sum_{i=1}^{n}\log(1+\phi_{i,k}^{\top}\Lambda_{i,k}^{-1}\phi_{i,k})$ instead) does not result in a substantial difference in computed trajectories, while resulting in more challenging numerical resolution of the resulting information gathering problem.
The information gathering objective (9) guides data collection towards regions of the feature space that enable rapid adaptation. As shown in Figure 3, the procedure of generating information gathering trajectories and adapting the model on them can be iterated as many times as needed to lower estimates of uncertainty. In our experiments, we found that adaptation on only one or two information gathering trajectories is sufficient for performing reliable drifting maneuvers, as shown in the next section.

### v2

@@ Abstract
Combining data-driven models that adapt online and model predictive control (MPC) has enabled effective control of nonlinear systems. However, when deployed on unstable systems, online adaptation may not be fast enough to ensure reliable simultaneous learning and control. For example, a controller on a vehicle executing highly dynamic maneuvers—such as drifting to avoid an obstacle—may push the vehicle’s tires to their friction limits, destabilizing the vehicle and allowing modeling errors to quickly compound and cause a loss of control. To address this challenge, we present an active information gathering framework for identifying vehicle dynamics as quickly as possible. We propose an expressive vehicle dynamics model that leverages Bayesian last-layer meta-learning to enable rapid online adaptation. The model’s uncertainty estimates are used to guide informative data collection and quickly improve the model prior to deployment. Dynamic drifting experiments on a Toyota Supra show that (i) the framework enables reliable control of a vehicle at the edge of stability, (ii) online adaptation alone may not suffice for zero-shot control and can lead to undesirable transient errors or spin-outs, and (iii) active data collection helps achieve reliable performance.
@@ I Introduction
Controlling unstable nonlinear systems remains challenging [1]. For example, driving a vehicle through dynamic and unstable drifting maneuvers is a difficult task where slight decision-making mistakes can lead to spin outs or crashes. In recent years, advanced control methods have been developed to give autonomous vehicles the ability to leverage their full handling potential [2, 3, 4, 5, 6], with possible applications to the future design of advanced driver-assistance systems and autonomous driving safety systems.
In this setting, data-driven model predictive control (MPC) methods have gained wide popularity [7] due to their high performance, capability to use expressive learned models, and ability to account for constraints such as actuator limits and obstacle avoidance. In particular, data-driven models that adapt using online data can reduce model errors and sim-to-real discrepancies at deployment time [8, 9, 10, 11]. However, simultaneously controlling and learning unstable systems is notoriously difficult. For instance (see Sec. V), successfully executing a drifting maneuver requires a sufficiently accurate initial model to precisely initiate the drift, and online adaptation alone may not be sufficient to execute the maneuver.
Autonomous driving at the limits presents a challenge for data-driven MPC with online adaptation due to its fast-paced nature and the inherent instability of the executed maneuvers. These difficulties raise the following questions: How can we effectively identify uncertain nonlinear dynamics to ensure that subsequent control is reliable? How should informative data be collected efficiently and safely to minimize time-intensive and potentially expensive testing on hardware?
Figure 1: We propose an active information gathering framework that guides data collection to quickly learn a model in a controlled (i.e., safe) setting. Adapting on this information-rich data enables reliable control in dynamic drifting trajectories. In contrast, directly using the model without adapting prior to deployment may lead to unreliable control performance, such as spinning out.
@@
  • We present a data-driven dynamics model tailored for driving at the limits of handling. This model combines prior physics knowledge with learned expressive neural network features. By adapting only the last layer, this data-driven model enables rapid online adaptation, significantly improving upon non-adaptive models in controlling unstable systems where small initial errors can quickly amplify.
  • To enable reliable deployment of the learned model in an MPC framework, we propose a pre-deployment active learning approach to identify an accurate model as fast as possible while satisfying constraints during data collection.
  • We extensively validate the proposed approach on a Toyota Supra in two different dynamic drifting maneuvers. In particular, results show that online adaptation of the meta-trained model alone may be insufficient to initiate a drift reliably and can lead to large transients in tracking error. On the other hand, the proposed active information gathering approach enables rapid identification of the system’s dynamics and yields improved tracking performance.
From a vehicle control perspective, our results demonstrate that executing drifting maneuvers is possible by adapting a model on informative non-drifting trajectories. Specifically, information-rich trajectories do not necessarily involve drifting, and adapting on such data significantly improves the execution of challenging drifting maneuvers.
@@ II Related Work
Data-driven models are increasingly popular for control due to their ability to leverage expressive models and adapt them online given measurements from the system, see [7] and [12] for recent reviews. For example, model predictive control (MPC) approaches leveraging neural network models have been widely used in recent years for dynamic driving at the limits of handling [9, 10, 11, 4]. These approaches adapt models online with the implicit assumption that the model will adapt fast enough during the control task to yield satisfactory performance, which may be difficult for systems where small initial errors quickly amplify.
Online learning and control algorithms that actively seek informative data to improve the model [13, 14, 15, 16, 17, 18, 19, 20] often result in improved control performance compared to passively adapting a model while performing a task, and allow enforcing safety constraints during the data collection process [15, 20]. Leveraging such active learning approaches for controlling unstable nonlinear systems remains challenging. Indeed, the tradeoffs between seeking informative data (exploration) and performing the task (exploitation) are well-known; unstable systems are particularly difficult to simultaneously learn and control [21, 22]. These tradeoffs motivate the design of a framework where active learning is done prior to deployment in a controlled setting to learn a model for control with minimal experimentation.
Autonomous driving at the limits of handling exemplifies these tradeoffs. In this application, the vehicle may operate with saturated tire forces close to the road boundaries, and it is desirable to identify the vehicle dynamics as accurately as possible prior to deployment. Previous approaches to driving at the limits leverage MPC with physics-based [5, 6] and neural network-based [9, 4, 11] dynamics models. Recently, [10] introduced a neural SDE dynamics model for MPC where the parameters of the SDE are sampled from a diffusion model conditioned on online observations, enabling passive online adaptation during task execution. Our proposed active learning framework complements these works, enabling the identification of uncertain dynamics as fast as possible prior to deployment in model-based control.
Active learning approaches to nonlinear system identification include methods that maximize information gain [16, 20, 23, 24]. Our approach relies on similar information-theoretic ideas to identify uncertain dynamics as quickly as possible. We leverage a Bayesian modeling approach [25] to last-layer meta-learning [26, 27, 25, 28, 29] that is conducive to information gathering [20]. Compared to driving at the limits of handling, experiments in [16, 20, 24] involve simulated or slow-paced robotic systems that do not face the same challenges of controlling and learning unstable dynamics. [23] also study active exploration for off-road driving, but use probabilistic ensemble neural networks that are computationally expensive compared to last-layer meta-trained neural network models, and penalize the vehicle sideslip which prevents executing drifting maneuvers. Additionally, as we show in Sec. V, online adaptation alone may not suffice to reliably execute drifting maneuvers, even using a meta-learned model capable of rapid adaptation.
@@ III Bayesian vehicle dynamics model
Central to our approach is a vehicle dynamics model that is expressive, captures uncertainty, and enables rapid online adaptation. To this end, we combine a physics-based nominal bike model with a Bayesian last-layer meta-learning model. The former encodes prior knowledge about the vehicle, while the latter captures model mismatch and enables rapid adaptation over expressive neural network features. The vehicle state is expressed in curvilinear coordinates [5, 11], see Fig. 2. As in [5, 11], we do not use brake inputs in this work. The states and control inputs of the vehicle are
@@
where $(n,m)=(6,2)$ are the state and control dimensions, $r$ is the yaw rate, $v$ is the total planar velocity, $\beta$ is the sideslip, $\omega_{r}$ is the rear wheelspeed, $e$ and $\Delta\varphi$ are the lateral and angle deviations to the reference trajectory, $\delta$ is the steering angle, and $\tau$ is the engine torque, respectively.
@@
where $k\in\mathbb{N}$ and $\epsilon_{k}\sim\mathcal{N}(0,\text{diag}(\sigma_{1}^{2},\dots,\sigma_{n}^{2}))$ are Gaussian-distributed disturbances that are independent over time $k$ and state dimensions. $\xi$ represents unknown parameters of the dynamics (e.g., corresponding to unmodeled phenomena) and $\tilde{u}_{k}=(u_{k},\dot{u}_{k})$ includes the control input rate $\dot{u}_{k}=(\dot{\delta}_{k},\dot{\tau}_{k})$ at time $k$. Including $\dot{u}_{k}$ in $\tilde{u}_{k}$ and (2) accounts for the fact that MPC plans $(u_{0},u_{1},\dots,u_{N})$ are executed by a low-level controller via linear interpolation on the control horizon. That is, $u(k\Delta t+\tau)=u_{k}+\tau\dot{u}_{k}$ with a time step $\Delta t$ and $\dot{u}_{k}=(u_{k+1}-u_{k})/\Delta t$ for $\tau\in[0,\Delta t)$.
@@
$x_{k+1}=h(x_{k},u_{k})+g(z_{k})+\epsilon_{k},$ (3)
where $h$ is the $\Delta t$-time step Euler integration of a standard bike model (see Appendix A for details). The residual $g$ in (3) is modeled by $g(z)=(g_{1},\dots,g_{n})(z)$ with inputs
@@
and each dimension $i$ of $g(\cdot)$ is linear in the learned features $\phi_{i}(z)\in^{d}$ for some parameters $\theta_{i}\in^{d}$, that is
@@
We omit the variables $e$ and $\Delta\varphi$ since they do not affect the dynamics of the other state variables, see [5]. That is, given the state in (1), we let $(g_{5},g_{6})(z)=0$ for all $z$.
The features $\phi_{i}(\cdot)$ are neural networks that are learned jointly with the parameters $\theta_{i}$ via last-layer Bayesian meta-learning [25, 30, 20] to yield an expressive model capable of rapid adaptation with uncertainty estimates. Specifically, viewing $g_{i}(z_{k})$ as a neural network, the last-layer parameter $\theta_{i}$ are parameterized using a Gaussian distribution
@@
with mean parameters $\bar{\theta}_{i}$ and positive definite precision matrices $\Lambda_{i}/\sigma_{i}^{2}$. Following [20], we model independent Gaussian distributions of each $\theta_{i}$, such that the joint distribution for the last layer parameters $(\theta_{1},\dots,\theta_{n})$ are Gaussian distributed.
Computing the posterior: The linear structure of $g_{i}(z)$ enables the computation of the last-layer parameters $\theta_{i}$ via Bayesian linear regression. Specifically, given prior distributions $\mathcal{N}(\bar{\theta}_{i,0},\sigma_{i}^{2}\Lambda_{i,0}^{-1})$ for $\theta_{i}$, and a state-control trajectory $\mathcal{D}_{T}=\{(z_{k},x_{k+1})\}_{k=0}^{T}$, the posterior $\theta_{i,T}|\mathcal{D}_{T}$ is Gaussian and can be computed recursively from $Q_{i,0}=\Lambda_{i,0}\bar{\theta}_{i,0}$ by
@@
With the proposed model, the mismatch between the true dynamics $f$ and the nominal model $h$ through $g$ is learned such that the one-step-ahead predictions are Gaussian by design, facilitating both meta-learning and information gathering.
Offline Bayesian meta-training: The prior parameters $(\bar{\theta}_{i,0},\Lambda_{i,0}^{-1})$, features $\phi_{i}$, and noise covariance scalings, $\sigma_{i}^{2}$, are pre-trained offline to obtain an expressive model that can be adapted online using the update rule (6). Specifically, given a dataset of $J$ trajectories corresponding to different parameters $\xi$ of the true system (2) and a meta-training horizon $T$, we minimize the negative posterior log-likelihood objective
@@
As a dataset, we train on a mix of successful and unsuccessful manual and autonomous drifting data totaling approximately 19 minutes of driving. This data corresponds to $35$ different trajectories, i.e., realizations of $f_{\xi}$, and contains a total of $11382$ data points. Following [11], we additionally augment our dataset with a mirrored set of data by negating the signs of the yaw rate, sideslip, and steering angle to simulate turning in the opposite direction. This doubles our dataset size.
For our architecture, we use a $\tanh$-feedforward neural network for each $\phi_{i}$ with two hidden layers of width $128$, and the dimension of each $\theta_{i}$ is $16$. Following [20], we share weights and biases across each $\phi_{i}$ except for their last layer weights and biases $(W_{\text{out}}^{i},b_{\text{out}}^{i})\in^{16\times 128}\times^{16}$. In total, there are $27012$ trainable parameters for the neural network, prior parameters, $(\bar{\theta}_{i,0},\Lambda_{i,0}^{-1})$, and noise covariance scalings, $\sigma_{i}^{2}$. We use a meta-training horizon $T=10$ and train our model using Adam with an initial learning rate of $10^{-3}$ and an exponential decay rate of $0.9975$ over $1000$ epochs.
@@ IV Information Gathering MPC Framework
We present an active information gathering framework, shown in Fig. 3, to identify nonlinear dynamics as fast as possible for reliable subsequent deployment within MPC. First, we meta-train a Bayesian vehicle model offline (Sec. III) using a dataset of vehicle trajectories. Second, we leverage the model’s expressive feature space $\phi$ and uncertainty representation encoded by $(\bar{\theta},\Lambda^{-1})$ to safely collect informative data and refine the model. After this information-guided adaptation, we leverage the refined vehicle model within an MPC framework to control the vehicle in challenging driving scenarios such as drifting. Next, we formulate the optimal control problems used for MPC and information gathering.
@@
We define an MPC formulation that enables tracking dynamic trajectories such as drifting maneuvers parameterized by a reference trajectory $x_{\mathrm{ref}}$. To favor smooth control inputs, we penalize fast changes in the control inputs and define
@@
$f_{\bar{\theta}}(x_{k},u_{k},u_{k+1})=h(x_{k},u_{k})+g(z_{k};\bar{\theta}),$
where $g_{i}(z_{k};\bar{\theta})=\bar{\theta}_{i}^{\top}\phi_{i}(z_{k})$ for each $i=1,\dots,n$.
Directly deploying an MPC controller that recursively solves OCP using the prior model parameters $\bar{\theta}_{i,0}$ after offline meta-training may lead to poor performance in practice, even if one adapts the model online during the task, see Sec. V. These observations motivate an active data collection approach to refine the model prior to deployment in challenging applications. Since data collection on a vehicle may be time-consuming and expensive, this raises the question: What data should we collect to best identify the unknown dynamics?
The key insight in answering this question is that learning the model as fast as possible amounts to maximizing the information gain from future observations, i.e., to maximizing the mutual information between observations and the true system (see [31, Chapter 2] and [32, Chapter 8]). The mutual information between two random variables characterizes their mutual dependence and provides a quantitative measure for selecting data that is most informative in identifying the true dynamics (2). Thus, to steer the system towards regions of the state-input space with high model uncertainty, we define the information gathering objective
@@
where $\alpha>0$ weighs the information gain compared to the nominal MPC objective. The nominal costs $\ell_{k}$ encourage smooth control inputs, and the constraints (8b)-(8e) ensure that the vehicle remains on the track during data collection.
The information gathering objective (9) is derived in [20] by leveraging the last-layer uncertainty representation of the model, $g$, which yields a closed-form expression of the one-step mutual information between an observation and the dynamics. Similar objectives are used for active learning in [33, 34, 16]. As discussed in [20], $\ell_{\mathrm{info}}$ is an approximation of the true total information gain over a trajectory, which is not exactly the sum of the expected information gains per timestep. We found that incorporating the linear regression updates (6) in the objective (by defining the cost ~$\tilde{\ell}_{\text{info}}=\frac{1}{2}\sum_{k=0}^{N-1}\sum_{i=1}^{n}\log(1+\phi_{i,k}^{\top}\Lambda_{i,k}^{-1}\phi_{i,k})$ instead) does not result in a substantial difference in computed trajectories, while resulting in a more challenging numerical resolution of the resulting information gathering problem.
The information gathering objective (9) guides data collection toward regions of the feature space that enable rapid adaptation. As shown in Fig. 3, the procedure of generating information gathering trajectories and adapting the model on them can be iterated as many times as needed to lower estimates of uncertainty. In our experiments, we found that adaptation on only one or two information gathering trajectories is sufficient for performing reliable drifting maneuvers, as shown in the next section.